hartnn (hartnn):

can anyone help me with Integration of (x^2)*(exp(-a*x)) from 0 to infinity

OpenStudy (unklerhaukus):

\[\int\limits_0^\infty x^2e^{-ax}\text dx\] you will have to use integration by parts twice \(\int u\text dv=\left. uv\right|-\int v\text du\) for the first time ,\[u=x^2\qquad\qquad \text dv=e^{-ax}\]\[\text du=2x\text dx\qquad\qquad v=\frac{e^{-ax}}{-a}\], dont forget the limits

OpenStudy (unklerhaukus):

* \(\text dv=e^{-ax}\text dx\)

hartnn (hartnn):

that i know,but is the final ans infinity or some function in only a....if latter,what is that function?

OpenStudy (unklerhaukus):

the final result is a function of a

hartnn (hartnn):

i m asking that function of a...i m getting infinity,which is clearly wrong

OpenStudy (lgbasallote):

is it just me or does this thing look a lot like laplace formulas :/

OpenStudy (unklerhaukus):

\[\int\limits_0^\infty x^2e^{-ax}\text dx=\left.x^2\frac{e^{-ax}}{-a}\right|_0^∞-\int\limits_0^∞2x\frac{e^{-ax}}{-a}\text dx\] \[=2\int\limits_0^∞x\frac{e^{-ax}}{a}\text dx\] now use integration by parts again \(\int p\text dq=\left. pq\right|-\int q\text dp\) \[p=x\qquad\qquad\text dq=\frac{e^{-ax}}{a}\text dx\]\[\text dp=\text dx\qquad\qquad q= \frac{e^{-ax}}{-a}\]

OpenStudy (unklerhaukus):

\[\qquad\qquad\qquad * q=\frac{e^{-ax}}{-a^2}\] \[2\int\limits_0^∞x\frac{e^{-ax}}{a}\text dx=2\left[\left.x\frac{e^{-ax}}{-a^2}\right|_0^∞-\int\frac{e^{-ax}}{-a^2}\text dx\right]\] \[\qquad=2\left[\int\frac{e^{-ax}}{a^2}\text dx\right]\]\[=2\left.\frac{e^{-ax}}{a^3}\right|_0^∞\]\[=\frac{2}{a^3}\]

OpenStudy (unklerhaukus):

BLAM!

OpenStudy (unklerhaukus):

any steps you not sure on, @hartnn?

OpenStudy (unklerhaukus):

note \[e^{-∞} =0\]

hartnn (hartnn):

actually i m not able to read/understand your response....the two responses before blam...

hartnn (hartnn):

actually i m not able to read/understand your response....the two responses before blam...

OpenStudy (unklerhaukus):

all i have done is to integrate by parts twice

OpenStudy (unklerhaukus):

can you be more specific ?

hartnn (hartnn):

got it,thanks so much :)

OpenStudy (unklerhaukus):

it all makes sense now/?

hartnn (hartnn):

but x^2*e^-ax when we put infinity,isn't it infinity*0???

hartnn (hartnn):

and not 0

OpenStudy (unklerhaukus):

\[\left.x^2\frac{e^{-ax}}{-a}\right|_0^∞=\left(\infty^2\frac{e^{-a\times\infty}}{-a}\right)-\left(0^2\frac{e^{-a\times0}}{-a}\right)\]\[\qquad\qquad=\left(\infty\times0\right)-\left(0\times0\right)\]\[\qquad=0-0\]\[=0\]

OpenStudy (unklerhaukus):

@lgbasallote is right, the easy way {if you know how}, is to recognise the laplace form \[\int\limits_0^\infty x^2e^{-ax}\text dx=\mathcal L\{x^2\}(a)=\frac{\Gamma(2+ 1)}{a^{2+1}}=\frac{2!}{a^3}=\frac2{a^3}\]